OpenStudy (precal):

Is there anyway to integrate y=sinx and y=sin(2x) 0 to pi by hand

OpenStudy (precal):

I have done the problem using my graphing calculator and double check in wolfram. my solution is 2.50 I broke it up into 2 parts \[\int\limits_{0}^{1.0471976}\sin(2x)-\sin(x) dx =.25\] \[\int\limits_{1.0471976}^{\pi}\sin(x)-\sin(2x) dx =2.25\] total is 2.5

OpenStudy (precal):

Is there anyway to do it by hand? My professor does not allow calculators but we are using a textbook that requires them.

OpenStudy (experimentx):

you know the standard values of sin(pi/2) and sin(pi)

OpenStudy (precal):

yes, sin(pi) =0 sin(pi/2)=1

OpenStudy (experimentx):

where did you break?? you should have broken down at pi/2

OpenStudy (precal):

I did not, I just stated the above trig values

OpenStudy (experimentx):

integrate from 0 to pi/2 sin(x) dx and integrate form pi/2 to pi sin(x) dx

OpenStudy (precal):

But my point of intersection is x=1.0471976

OpenStudy (precal):

pi/2 is not 1.0471976

OpenStudy (experimentx):

I got the same result ... I think wolfram is doing something/ http://www.wolframalpha.com/input/?i=integrate+from+0+to+1.0471976+sin%28x%29+%2B+integrate+from+1.0471976+to+pi+sin%28x%29

OpenStudy (precal):

you forgot sin(2x)

OpenStudy (experimentx):

and that value you mentioned is nearly pi/3 http://www.wolframalpha.com/input/?i=pi%2F1.0471976

OpenStudy (experimentx):

oops .. sorry.

OpenStudy (precal):

yes you are correct, it is pi/3

OpenStudy (precal):

How do I solve it by hand?

OpenStudy (precal):

sin(x)=sin(2x)

OpenStudy (experimentx):

pi/3 has standard values ...and all it's multiples. I guess express it as surds.

OpenStudy (experimentx):

what exactly are you trying to do??

OpenStudy (precal):

wolf can't show steps

OpenStudy (precal):

I have to do this problem without using any technology

OpenStudy (experimentx):

yes ... but is it just integration??

OpenStudy (precal):

yes

OpenStudy (precal):

\[\int\limits_{0}^{\frac{\pi}{3}}\sin(2x)-\sin(x) dx + \int\limits_{\frac{\pi}{3}}^{\pi}\sin(x)-\sin(2x) dx\]

OpenStudy (experimentx):

why are you breaking it in parts??

OpenStudy (precal):